can you solve this using substitution?
-2x+y=-11
4x+4y=4

Answer by Alan3354(69443)

can you solve this using substitution?
-2x+y=-11
4x+4y=4
-------------
Yes, I can.
Can you?
==============
-2x+y=-11
4x+4y=4 --> x+y = 1
----------------------------------
-2x+y=-11
x + y = 1
----------------
Solve for y in the 2nd eqn.
Sub for y in the 1st eqn.
----------
Or vice versa.